OpenStudy (anonymous):

if sinusoidal functions is not even then its derivative is even. true or false?

OpenStudy (anonymous):

True. The derivative of any odd function is always even.

OpenStudy (bennettlafayette):

And for the proof of such we can use the chain rule and definition of an odd function: f is odd if the following is true: f(x) = - f(-x) let z = -x dz/dx = -1 [equivalently z' = -1] Make a z substitution in the definition of the odd function f(x) f(x) = - f(z) differentiate using chain rule: df(x)/dx =-[ df(z)/dz ] * [dz/dx] {equivalently, f '(x) = - f '(z) z' } df(x)/dx = -[df(z)/dz] * (-1) so, df(x)/dx = df(z)/dz f '(x) = f '(z) substitute for z = -x f '(x) = f '(-x) satisfies the definition of an even function and since we learned in lecture that f ' is indeed a function it is also an even function. So, derivatives of odd functions are even.

OpenStudy (arnavguddu):

given f(x)=-f(-x) =>odd diff both side f'(x)=-(-f'(-x)) or f'(x)=f'(-x) => even so odd sinusoid has even derv

OpenStudy (arnavguddu):

and vice-versa